### v1

# Ready Jurist One: Benchmarking Language Agents for Legal Intelligence in Dynamic Environments

Abstract
The gap between static benchmarks and the dynamic nature of real-world legal practice poses a key barrier to advancing legal intelligence. To this end, we introduce J1-E NVS , the first interactive and dynamic legal environment tailored for LLM-based agents. Guided by legal experts, it comprises six representative scenarios from Chinese legal practices across three levels of environmental complexity. We further introduce J1-E VAL , a fine-grained evaluation framework, designed to assess both task performance and procedural compliance across varying levels of legal proficiency. Extensive experiments on 17 LLM agents reveal that, while many models demonstrate solid legal knowledge, they struggle with procedural execution in dynamic settings. Even the SOTA model, GPT-4o, falls short of 60% overall performance. These findings highlight persistent challenges in achieving dynamic legal intelligence and offer valuable insights to guide future research.
Figure 1: An illustration of six legal environments in J1-E NVS , categorized into three levels by environmental complexity: Level I covers legal knowledge questioning and consultations; Level II includes complaint drafting and defence drafting; Level III involves civil court and criminal court.
1 Introduction
Legal Benchmarks Multi-turn? Dynamic? Procedural Process? Type Evaluation
LexGLUE (ACL 2022) Chalkidis et al. (2021) NLP Task Rule
LegalBench (NeurIPS 2023) Guha et al. (2023) NLP Task Rule
LawBench (EMNLP 2024) Fei et al. (2024) NLP Task Rule
Law-Eval (DASSFA 2024) Yue et al. (2024) NLP Task Rule
Law-Eval (DASSFA 2024) Yue et al. (2024) NLP Task Rule & LLM
LAiW (COLING 2024) Dai et al. (2023) NLP Task Rule & Human
J1-E VAL (Ours) Scenario Rule & LLM
Table 1: Comparison between our J1-E VAL and existing legal benchmarks. Our benchmark is scenario-based, dynamic, and interactive, aligning better with real-world legal practices.
Advances in Large Language Models (LLMs) have shown remarkable capabilities across a range of traditional legal NLP tasks, such as legal information extraction Yao et al. (2022); Shen et al. (2024), judgment prediction Huang et al. (2021); Song and Wei (2021). More recently, LLM-based agents are expected to accomplish complex tasks within dynamic environments through sustained, interactive engagement Xi et al. (2024); Mou et al. (2025); Yue et al. (2025a), thereby bringing substantial potential for expanding the scope of legal intelligence.
To study such interactive and dynamic legal intelligence, the first step is to establish a reliable evaluation system to identify the gap between current LLM-based and ideal legal agents. As shown in Table 1, existing legal benchmarks Li et al. (2024); Fei et al. (2024); Yue et al. (2024) predominantly adopt static and non-interactive paradigms, where direct predictions are based on predefined legal elements, such as multiple-choice questions and conventional legal task reconstructions. However, these methods struggle to capture the dynamic and professional nature of real-world legal practice. Realistic services involve multi-turn interactions among practitioners, as the public often lacks legal knowledge and requires step-by-step guidance from legal professionals. Moreover, this process also emphasizes the procedural legality. This misalignment between current methods and realistic situations limits the understanding of language agents’ capabilities in legal contexts.
Studies on LLM-based agents in legal contexts should focus on their ability to operate in dynamic environments, mirroring the real-world interplay of legal practitioners with evolving information and procedures. Ideal environment, e.g., court trials, relies on human participants with defined roles to ensure reliability, yet such setups are prohibitively costly and difficult to scale. Thus, developing such a system faces two key challenges: 1) Comprehensive scenarios. Real-world legal scenarios encompass diverse participants and procedural stages, e.g., advising the public, drafting documents, or issuing judgments. Sufficiently rich and representative scenarios are essential for uncovering agents’ legal sense across multiple dimensions. 2) Real-world simulation. Each scenario involves diverse roles with distinct behaviors, backgrounds, and legal perspectives, all requiring a realistic and consistent portrayal. Beyond individuals, interaction must be dynamic yet procedurally grounded.
To this end, we present J1-E NVS , an open-ended legal environment where agents engage in diverse legal scenarios, completing tasks through interaction with varied participants under procedural norms. Guided by legal experts, it curates six representative environments, categorized into three levels according to environmental complexity (role diversity, interaction demands, and procedural difficulty) as shown in Figure 1. (1) Level-I comprises knowledge questioning and legal consultation: the legal agent answers the public’s progressive questions on legal knowledge or specific cases. (2) Level-II involves complaint and defence drafting: the legal agent provides step-by-step guidance to litigants in compliance with procedural requirements, and systematically collects the necessary information to complete the drafting process. (3) Level-III consists of civil and criminal courts: within formal judicial processes, the legal agent facilitates interactions among parties, ensures procedural compliance, and produces legally valid judgments. Each environmental role is grounded in real-world legal contexts and exhibits a distinct behavioral style, ensuring diversity and authenticity.
Furthermore, we introduce J1-E VAL to evaluate agent capabilities within our constructed environments. We design task-specific and fine-grained metrics to assess task completion and procedural compliance, providing ground truth for each task and adopting rule-based or LLM-based methods. These fine-grained metrics give a deeper understanding of the skill sets required for legal task execution, and offer a realistic and reliable overview of agent capabilities, facilitating further exploration into the similarities and differences between LLM-based agents and human learning mechanisms.
We extensively evaluate 17 popular LLM-based agents, including proprietary models, open-source models, and legal-specific models. The results indicate that while current agents have internalized legal knowledge, they still struggle to follow procedural steps when performing tasks in dynamic environments. Even the SOTA model (GPT-4o) fails to surpass the average score of 60, highlighting the ongoing challenges in achieving legal intelligence.
Contributions: (A) We introduce the first interactive and dynamic legal environment for agents. Our J1-E NVS covers six of the most representative real-world scenarios in Chinese legal practice. (B) We introduce J1-E VAL to perform fine-grained evaluation of agents across different levels of legal proficiency, ranging from trainee to lawyer to judge. (C) Experimental results reveal gaps between diverse agents and real-world legal demands, offering valuable insights to guide future progress. (D) This work introduces a new paradigm for legal intelligence, shifting from static to dynamic. Beyond evaluation, it can further be extended for data generation and reinforcement learning training.
Figure 2: Environment construction pipeline. (A) Role Agent Setting: We leverage real-world legal sources and personality theories to construct different role agents. (B) Multi-level Environment Construction: We combine the various roles in specific procedures and relationships to form legal environments.
2 J1-ENVS: Interactive Legal Environments
This section elaborates on J1-E NVS , which encompasses six representative scenarios derived from Chinese judicial practice. We detail its construction pipeline from individuals to communities as shown in Figure 2: (A) Role Agent Setting2.1) leverages real-world legal sources and personality theories to construct different role agents; (B) Multi-level Environment Construction2.2) combines the various roles in specific procedures and relationships to form legal environments.
2.1 Role Agent Setting
Legal environments encompass a variety of roles, including legal professionals with domain expertise and non-legal participants with diverse needs, behaviors, and backgrounds. Due to limited expertise, generic LLM-driven simulators are prone to substantial legal illusions. Thus, realistic role agent settings face two challenges: diversity of non-legal participants and legal consistency among roles. We address this by modeling individuals with the Big Five personality traits and assigning consistent legal logic to agents in real-world legal scenarios.
Setup with Real-world Legal Sources. To simulate realistic scenarios, we assign different legal elements of the same event to various roles, drawing from two sources: (1) Chinese judgment documents. For scenarios in Level II & III, Chinese court judgments encompass all the essential components of a case’s entire procedural process, e.g., parties’ details, legal disputes, facts, rulings, and laws, which align perfectly with our requirements. Specifically, we collect civil and criminal cases from judgment documents and then employ LLM to extract key elements, such as plaintiff information, claims, case details, and relevant legal provisions. (2) Chinese legal articles. For scenarios in Level I focused on progressive questioning logic and authenticity, we collect expert-drafted legal articles from the HUALV website and the Civil Trial Practice Q&A literature. As these articles are structured through stepwise questions or thematic headings, we prompt LLM to leverage this inherent structure to generate coherent topic lists and corresponding answers. Topics are classified as binary (yes/no) and open-ended questions. Details are provided in Appendix B.1.
Setup with Personality Theory. Behavioral diversity enhances the model to represent complex characters, further aligning real-world behavior. Guided by social personality theory Sun et al. (2024), Big-5 Personality Traits are mapped to behavioral styles of non-legal roles (public, the plaintiff, and the defendant). Each trait is assigned one of three levels (high, medium, or low) and GPT-4o is prompted to generate corresponding behavioral descriptions based on the role’s profile or background. Details are provided in Appendix B.2.
Role Agents Configuration. As shown in Figure 2, different elements are assigned to three levels of roles. (1) Type I: General Public represents persons in various occupations, such as workers and farmers, and typically lack legal knowledge and thus seek knowledge-based (Env.1) or case-specific (Env.2) consultations. They are assigned topic lists, background, and behavioral styles. (2) Type II: Specific Characters have more explicit litigation needs. They are assigned individual profiles, including personal attributes (e.g., names, genders, and addresses) and case-related details (e.g., claims, statements of defence, and evidence). Besides, they also configure behavioral styles. (3) Type III: Multi-specific Characters hold specific responsibilities within courtroom proceedings and are assigned real-world legal information corresponding to their roles. For example, the Lawyers in the civil court possess detailed case information extracted from civil judgment documents, e.g., their client information, case details, claims & defence. In criminal court scenarios sourced from criminal judgment documents, the Procurator agent is assigned to charges and opinions, while the Defendant agent is assigned case details and a stated stance toward the charges. See Appendix B.3 for details.
2.2 Multi-level Environment Construction
Building upon curated role sets, we construct six representative environments grounded in Chinese legal practice, spanning the full spectrum of legal engagement from initial inquiry to formal adjudication. These environments are classified into three levels by role diversity, interaction complexity, and procedural depth.
Level I: Knowledge Questioning (K Q ) and Legal Consultation (L C ).
This level involves two participants, where the general public initiates questions, and the legal agent responds. It aims to simulate the response capabilities of the legal trainee in dynamic environments. (1) K Q . Through progressive dialogue, the legal agent resolves user confusion, moving from broad themes to specific issues. The public, assigned a topic list and occupation, continues the interaction until all topics are addressed. (2) L C . This presents a complex counseling setting grounded in a specific incident. Since legal agents lack information about the public, they actively ask for relevant details to provide better responses. Similar to KQ, the interaction continues until all question lists are resolved.
Level II: Complaint Drafting (C D ) and Defence Drafting (D D ).
This level involves two participants: an individual with a specific legal need, and a legal agent who guides the interaction to progressively collect the required information and ultimately address the litigant’s request. It aims to simulate the process by which a practicing lawyer independently initiates and completes tasks based on a given agenda. (1) C D . The legal agent gathers information from the plaintiff and formulates a complete complaint as the final objective. Guided by a predefined format, the agent autonomously plans the required content through interaction. The scenario ends once the complaint is successfully generated. (2) D D . In response to the defendant’s need for a statement of defence, the legal agent engages in in-depth dialogue with the defendant, drawing on the complaint to organize arguments and ultimately generate the defence document. The end of conditions is similar to complaint drafting.
Level III: Civil Court (C I ) and Criminal Court (C R ).
This level involves multiple participants and is governed by strict court procedural norms. (1) C I . The Civil Court addresses disputes concerning rights and obligations between individuals or organizations, involving three participants: the plaintiff, the defendant, and the judge. It includes five stages: court preparation, court investigation, court debate, court mitigation, and court decision. The judge agent oversees the process by following the procedural framework and interacting with both parties’ attorneys. In the final stage, the agent renders a judgment in response to the plaintiff’s claims. (1) C R . The Criminal Court centers on adjudicating criminal charges against individuals or groups, involving four participants: the defendant, the defendant’s lawyer, the procurator, and the judge. It includes three stages, court preparation, court investigation, and decision. Compared to civil court, legal agents in criminal proceedings interact with a broader range of participants, eliciting statements from the defendant, requesting evidence from the prosecutor, and working toward the establishment of criminal liability.
3 J1-Eval: Holistic Legal Agent Evaluation
Figure 3: Distribution of legal attributes for six environments in J1-E VAL , showing a wide range of coverage.
3.1 Data Details
Following the details in Sec.2.1, we collected legal articles and documents from the year 2025 to ensure the timeliness. After systematic processing, J1-E VAL sets out 508 distinct environments, including 160 at Level I, 186 at Level II, and 192 at Level III. The detailed statistics are presented in Table 4 in Appendix. Moreover, J1-E VAL encompasses a diverse and nuanced set of legal attributes aligned with real-world legal practice. As illustrated in Figure 3, K Q and L C involve different entities and procedural topics.C D , D D and C I address various types of civil rights disputes. C R covers different categories of criminal offenses.
3.2 Evaluation Methods
Considering the distinct characteristics of each legal task, we design tailored metrics accordingly, which enables a clearer understanding of the legal agent’s fine-grained performance. The legal agent needs to accomplish the corresponding legal tasks in dynamic environments. For each task, we adopt either rule-based or LLM-based automatic evaluation methods with explicit ground truth references. Each metric adopts a score range in [lower bound–upper bound] form. Due to page limitations, we provide a brief overview, see Appendix D for full details.
(1) Level I.
To evaluate progressive questioning and responses, we implement an LLM-based check mechanism to label each interaction round with its corresponding topic, facilitating <question–answer> matching. The two setting types provide a comprehensive assessment: for binary questions, outputs are summarized as "yes" or "no," while for non-binary questions, original responses are used for evaluation.
  • Binary accuracy (B IN ) [0, 0.5, 1] measures accuracy in binary questions by matching method: 1 for correct, 0 for incorrect, and 0.5 for conditionally qualified answers.
  • Non-binary score (NB IN ) [0-10] evaluates the response based on ground truth in terms of legal points, logic, and factual accuracy. We use GPT-4o to implement it.
(2) Level II.
This stage involves the task of document generation, where the quality reflects the completeness of litigant information and evidence, the understanding of claims, and the appropriate use of document templates. Two metrics are used:
  • Format-following score (F OR ) [0-1] evaluates how closely the format of the generated legal document matches the reference, based on key components (e.g., plaintiff, claims, evidence), their order, and correct labeling.
  • Document score (D OC ) [0,1 or 0-10] measures the average accuracy of each component of document: personal information (Plaintiff or defendant) is scored using exact match (0 or 1), while other components are evaluated by GPT-4o based on their consistency with the reference answer.
(3) Level III.
This stage evaluates the quality of the final judgment and procedural compliance. Given the differences between criminal and civil judgments, we adopt distinct methods. The evaluation metrics are as follows:
  • Procedural-following score (P FS ) [0-1] evaluates completeness of procedural stage. The court contains several stages, each comprising a set of mandatory actions defined based on the legal professional’s setting. A stage is deemed complete only when all presetting actions within that stage are correctly matched.
  • Judgment score (J UD ) [0-1 or 1-10] evaluates the judgment quality. For civil court, LLM evaluates the judgment against the ground truth regarding legal points, logicality, and factual accuracy. For criminal court, we use matching to calculate the crime (C RI ) accuracy. Additionally, following Chen et al. (2019), log distance is applied to calculate the deviation span for verdict (V ER ): fines and sentences. Lower V ER indicates higher accuracy.
  • Reason score (R EA ) [1-10] evaluates the reason based on ground truth in terms of legal points, logic, and factual accuracy .
  • Law accuracy (L AW ) [0-1] measures law accuracy by matching based on ground truth.
Model Level-I Level-II Level-III
KQ LC CD DD CI CR
B IN N B IN B IN NB IN F OR D OC F OR D OC P FS J UD R EA L AW P FS J UD R EA L AW
C RI V ER
Multilingual LLMs (close source)
GPT-4o 69.70 62.73 55.85 45.05 58.92 89.86 45.38 84.79 64.22 31.18 53.12 14.12 29.17 92.75 7.65 63.19 41.45
Claude-3.7 68.41 60.69 52.42 49.57 68.82 89.23 38.71 68.76 59.69 26.88 46.45 17.83 25.36 89.86 14.28 73.04 41.81
Multilingual LLMs (open source)
Deepseek-v3 671B 67.11 60.71 54.50 49.65 84.73 63.60 57.20 76.25 28.85 14.30 19.57 10.83 16.39 81.16 30.76 60.29 33.44
Deepseek-r1 671B 72.46 62.78 57.19 48.28 69.89 46.44 36.34 47.89 14.42 7.63 9.68 5.13 2.45 7.25 97.88 6.09 4.08
Llama3.3-it. 70B 64.76 50.07 43.88 23.60 97.85 84.52 32.26 62.35 56.01 20.54 38.17 6.07 21.83 86.96 19.76 54.49 18.62
Qwen3-it. 32B 69.64 66.16 58.33 47.24 91.18 84.42 46.67 63.46 45.99 22.69 43.87 19.19 19.02 82.61 20.94 68.12 39.35
Gemma3-it 27B 64.09 49.50 54.10 26.01 96.34 84.83 37.85 68.82 52.05 21.83 43.87 7.12 28.71 52.17 17.04 37.68 17.60
Qwen3-it. 14B 67.70 62.84 56.18 44.33 63.87 89.15 36.56 62.38 1.40 1.40 1.83 1.09 31.88 85.51 16.96 65.07 47.04
Gemma3-it 12B 62.03 41.98 47.85 19.22 51.83 84.14 35.27 63.85 43.94 16.24 34.19 3.91 20.38 82.61 15.16 46.81 12.39
GLM4 9B 65.95 54.86 54.03 32.35 56.56 82.21 40.43 52.20 44.85 12.47 26.45 3.22 24.82 68.12 42.09 41.88 17.91
Qwen3-it. 8B 66.24 57.14 54.97 36.90 74.19 72.70 17.63 22.82 12.84 10.32 21.72 10.06 8.61 50.72 42.50 36.52 29.48
InternLM3-It. 8B 65.25 55.26 55.51 42.35 90.97 72.19 36.13 53.82 8.30 11.29 12.15 - 17.66 47.83 58.80 14.49 0.41
Ministral-it. 8B 59.60 37.60 51.48 14.53 60.00 2.84 34.19 13.99 25.23 7.42 16.88 0.54 2.17 5.80 96.06 3.48 1.09
Qwen2.5-it. 7B 67.21 54.92 52.60 31.28 89.46 81.41 47.10 59.71 23.40 1.61 10.22 3.49 21.65 55.07 46.90 36.23 26.85
Qwen3-it. 4B 62.17 55.90 50.67 37.15 51.40 83.68 33.12 51.45 26.74 13.87 23.01 8.99 15.58 65.22 38.32 48.70 26.99
Legal-specific LLMs
LawLLM 13B 56.25 47.25 44.92 20.79 12.90 2.85 - - - - - - - - 100.00 - -
Chatlaw2 7B 60.64 53.11 50.40 25.01 22.37 16.97 1.29 0.89 - 1.08 0.54 - - - 98.55 - -
Table 2: Performances among LLM-driven legal agents on J1-E NVS , where the bolded score indicates the best performance, while the underlined score represents the second-best. Note that “-” denotes failure to accomplish the task, and lower V ER indicates higher performance.
4 Experiments
4.1 Experimental Settings
Agent Models.
We evaluate legal agents driven by various LLMs in two categories. (1) General multilingual LLMs: GPT-4o111gpt-40-2024-11-20 Achiam et al. (2023), Claude-3.7222claude-3-7-sonnet-20250219 Anthropic (2023), Deepseek-v3 Liu et al. (2024), Deepseek-r1 Guo et al. (2025), InternLM3-instruct 8B Cai et al. (2024), Qwen2.5-instruct 7B Yang et al. (2024), Qwen3-instruct 4B/8B/14B/32B Yang et al. (2025), ChatGLM4-chat 9B GLM et al. (2024), and Ministral-Instruct-2410 8B Jiang et al. (2023a). (2) Legal-specific LLMs: LawLLM 13B Yue et al. (2024), and Chatlaw2 7B Cui et al. (2023).
Implementation Detail.
We use Qwen3-Instruct 32B Yang et al. (2025) and GPT-4o Achiam et al. (2023) to drive different environments. Due to page limitations, details of the evaluation are provided in Appendix E.
4.2 Performance among Various Legal Agents
In this section, we use GPT-4o Achiam et al. (2023) to drive our environments, where LLM-base agents enter these to accomplish set legal tasks.
Overall Performance.
To assess overall performance, we compute the average scores across all environments except V ER , and rank the models by size. As shown in Figure 5, GPT-4o achieved the best performance, showing strong legal intelligence. Qwen3-Instruct-32B demonstrates performance beyond expectations, surpassing Deepseek series. Notably, although the legal-specific LLMs perform comparably to the GPT series on existing legal benchmarks Fei et al. (2024); Yue et al. (2024), they exhibit significantly weaker performance in our setting, falling behind even smaller models. This highlights a key limitation: despite possessing legal knowledge, the absence of interactive capability hampers their effectiveness in dynamic, real-world environments. Overall, legal intelligence tends to improve with increasing model size. However, even the SOTA model fails to surpass an average score of 60, underscoring the persistent challenges that LLM agents face in handling diverse and complex legal scenarios.
Figure 4: Overall performance ranking across different LLM agent sizes.
Figure 5: Completion rate of different stages in civil court (Env.5) and criminal court(Env.6).
Performance at Level I & II.
Table 2 reports our fine-grained metrics across levels. At Level I, both general and legal-specific models perform relatively well on K Q , indicating a solid understanding of legal knowledge. However, their performance drops on the more interactive L C task, which requires proactive engagement. Notably, even legal-specific agents built on older base can match the performance of newer LLMs. Results indicate that current agents possess legal knowledge and can apply it effectively in basic legal consultations. At Level II, LLaMA-3.3 and Deepseek-v3 achieve the highest scores on F OR for C D and D D tasks, demonstrating strong template adherence in the long legal context. For document generation quality, GPT-4o performs the best. By contrast, legal-specific and smaller models, e.g., Ministral-Instruct and LawLLM, fail to complete the assigned tasks. Detailed results of the document (D OC ) components are shown in Table 6 and 7 of Appendix.
Performance at Level III.
Civil and criminal courts involve distinct procedures and judgment forms. The civil court comprises five stages, yet most models struggle to complete the full process (P FS ), particularly in legal-specific models, reasoning models, and smaller models. GPT-4o and Claude-3.7 perform well in P FS , R EA and L AW . The criminal court includes three stages and more participants, where closed-source models achieve similar leading performance. Additionally, Deepseek-r1’s poor performance in executing the process highlights the limitations of reasoning models in complex environments. In summary, effective reasoning and adherence to procedural protocols within long legal contexts are essential for achieving accurate legal judgments. Detailed results of judgment components are shown in Tables 8 and 9 of Appendix.
Performance of Procedural-Following Ability
To further analyze the performance of legal agents in the civil and criminal court, we measure the completion rate by computing how many samples successfully reach designated stages. As shown in Figure 5, legal agents struggle to complete necessary actions in the Preparation stage in both civil and criminal courts. In addition, despite fewer stages and actions, navigating through the criminal court poses greater challenges for legal agents, as the environment contains more characters that increase the complexity of interaction. These findings underscore the need for enhanced procedural learning for legal agents, as well as the importance of strengthening their communication and coordination with different characters.
4.3 Can LLMs drive J1-ENVS?
Existing works Fan et al. (2025); Mou et al. (2025) have shown the strong role-playing and community modeling abilities of LLMs. In this section, we use GPT-4o and Qwen3-Instruct-32B to drive the J1-E NVS separately, demonstrating the reliability.
Agent Env.GPT-4o Env.Qwen3-it. 32B
Evaluator(GPT-4o)
GPT-4o 94.22 95.49
Deepseek-v3 671B 90.97 94.71
Qwen3-it. 14B 89.97 94.99
Evaluator(Human)
GPT-4o 86.67 84.33
Deepseek-v3 671B 84.30 83.15
Qwen3-it. 14B 86.33 86.39
Table 3: Behavior consistency with various legal agents under Human and GPT-4o evaluation.
Analysis of behavioral consistency.
We analyze the behavioral consistency of environment roles during interactions with three legal agents. Both GPT-4o and human evaluators rate the consistency between each role’s profile and its behavior on a scale of 1 to 10, GPT-4o evaluated all samples, while human evaluators assessed 10% of the samples per scenario. The average scores are shown in the Table 3. Across interactions with various legal agents (GPT-4o, Deepseek-v3, and Qwen3-Instruct-14B), the ratings remain consistently high and stable. Furthermore, the scores for interactions with the same agent are comparable regardless of environments. This experiment validates the reliability and effectiveness of our environment, laying a solid foundation for assessing agents’ performance in interactive legal scenarios.
Figure 6: Overall performance of legal agents with J1-E NVS driven by GPT-4o or Qwen3-Instruct-32B.
Figure 7: The distribution of the difference between GPT-4o scores and human evaluation
Analysis with different J1-E NVS .
To further validate the stability, we employ a smaller, publicly accessible LLM (Qwen3-Instuct-32B) to drive our environment, and evaluate the overall performance of five legal agents, i.e, GPT-4o, Deepseek-V3, Qwen3-Instuct-14B, Gemma3-Instuct-14B and Chatlaw2. Figure 7 illustrates that: 1) all five agents successfully completed their respective tasks within the Qwen-based environment and achieved the same rankings as in the GPT-4o-based environment; 2) under the Qwen3-Instruct-32B environment, the performance of each agent is lower than that in the GPT-based setup, which is an acceptable outcome given the gap in model size and capability. Detailed results are shown in Table 5. Overall, the relative performance differences and rankings among agents remain consistent across various environments, further validating the robustness of our framework and highlighting its potential for deployment across diverse scenarios.
Analysis of the LLM evaluator in J1-E VAL .
We analyze the reliability of the LLM-based evaluation of various metrics, such as N BIN , D OC , J UD , and R EA . We select 10 cases per scenario. Human evaluators are asked to evaluate the performance of legal agents, and the difference between the human score and the corresponding GPT-4o score is plotted in Figure 7. We find that the scores provided by GPT-4o and human evaluators largely(over 76%) remain consistent. These results demonstrate the reliability and robustness of our evaluation framework, which enhances the validity of the findings of the legal agents.
5 Conclusion
This paper introduces J1-E NVS , an environment designed to simulate agents’ legal interactions in various legal scenarios. We further presents J1-E VAL to assess agents’ ability to operate effectively within dynamic environments. Distinct from prior benchmarks, our framework emphasizes interactivity and encompasses a broad range of realistic legal tasks. Experimental results demonstrate the stability and effectiveness of dynamic environments in J1-E NVS . By evaluating a wide range of existing general-purpose and legal-specific LLMs as legal agents, the study uncovers both notable advancements and persistent challenges faced by current agents. Our framework not only functions as an evaluation system but also serves as a reliable environmental toolkit for data generation or reinforcement learning training.
Limitations
Although this work presents the first comprehensive legal dynamics benchmark based on an LLM-driven framework, designed to closely simulate real-world legal practice scenarios, certain limitations remain in the current evaluation of agents. In particular, the benchmark focuses primarily on the procedural flow of legal tasks, while in real-world applications, agents may require more complex capabilities to complete tasks effectively, such as retrieving statutory provisions or consulting precedent databases. These aspects are not yet considered in the current framework. Future research is expected to build upon this foundation by incorporating such advanced functionalities, thereby further improving the realism and applicability.
Appendix A Related Work
Legal Intelligence.
Legal AI applies artificial intelligence to enhance or automate various tasks in the legal domain, such as legal information extraction Yao et al. [2022], Shen et al. [2024], judgment prediction Huang et al. [2021], and legal question answering Kien et al. [2020], Zhong et al. [2020]. Early approaches Zhong et al. [2020], Ji et al. [2021] relies on task-specific models and curated datasets for specific legal tasks. With LLMs’ strong generalizability Brown et al. [2020], they have shown notable capabilities in the legal domain. For example, Deng et al. Deng et al. [2023] introduced a structured prompting method to enhance judgment prediction. ChatLaw Cui et al. [2023] and LawLLM Yue et al. [2023, 2024] enhance multitasking capabilities through training on legal knowledge and instructions. Recent explorations Yue et al. [2025b], Chen et al. [2024] have employed LLM-driven agents to simulate legal scenarios. However, a key challenge remains: the lack of a unified framework for evaluating agent performance and guiding future progress.
Legal Benchmark.
Existing legal assessments of LLM are predominantly static, single-turn paradigms Fei et al. [2024], Li et al. [2024], Guha et al. [2023], Dai et al. [2023], Yue et al. [2024]. Specifically, Law-Eval Yue et al. [2024] evaluates the LLM using multiple-choice questions from law-related exams. LawBenchFei et al. [2024]and LegalBenchGuha et al. [2023] evaluate various dimensions of legal competence by consolidating existing legal tasks into an LLM-adapted format. Furthermore, LAiW Dai et al. [2023] follows the logic of legal practice of Syllogism by dividing the legal capabilities of LLMs into three levels. Inspired by Bloom’s taxonomy, LexEval Li et al. [2024] proposes a new taxonomy of legal cognitive abilities to organize different tasks. Existing benchmarks adopt a static framework centered on a singular dimension of task competence. Such an approach is inadequate for capturing the dynamic nature of interactions and fails to account for the complex, multi-skill demands of realistic scenarios. Consequently, they are insufficient for the emerging paradigm of agent-based evaluation.
LLM-based Simulation.
By mimicking human senses and vivid performance, LLMs can be leveraged to power role agents with diverse personalities and environmental contexts, ranging from individuals Shao et al. [2023], Argyle et al. [2023] to demographic groups Zhang et al. [2025], Jiang et al. [2023b]. These agents can coordinate, collaborate, exchange information, or even conduct debates, such as providing information through interaction, showing social intelligence Zhou et al. [2023], Mou et al. [2025]. Based on role agents, some works attempt to build more complex scenarios Stade et al. [2024], Fan et al. [2025]. For example, Generative Agents Park et al. [2023] construct a small town with 25 agents. AI hospital Fan et al. [2025] simulates dynamic medical interactions between a doctor and NPCs, including a patient, an examiner, and a chief physician. The powerful role-representation enabled by LLMs lays the foundation for constructing complex community simulations.
Appendix B Role Setting Details
B.1 Legal Sources Processing
We utilize two primary legal sources to construct the role-specific settings.
Chinese Judgment Documents. Judgment documents are official court records of judicial proceedings, outcomes, and factual details regarding involving parties. We collect both civil and criminal judgment documents, and employ an extraction model(GPT-4o) to identify and structure relevant legal elements. For civil judgment documents, we extract the following categories of legal elements.
  • Personal information includes two types. For individual-specific elements, we collect name, gender, ethnicity, birthdate, and residential address. For legal entities(corporations), personal attributes include the corporate name, address, and the identity of the legal representative. These elements are subsequently assigned to the roles of plaintiff and defendant in Complaint & Defence Drafting, as well as Civil Court.
  • Case information captures the procedural details of the case, including plaintiff claims, case details, the defendant’s defence against accusations, and evidences.
  • Court Information includes all the necessary elements for court rulings, containing the final judgment, court findings about the case, and laws applied by the court in the case.
For criminal judgment documents, the following legal elements are derived:
  • Personal information includes key attributes of the defendant, encompassing name, ethnicity, address, gender, birth date, occupation, education level, and custody status.
  • Plea captures the positions of the defendant and their lawyer on the charges, including statements and responses towards accusations.
  • Court information contains the final verdict, court findings, and legal provisions cited by the judge, resembling the information extracted from civil documents.
  • Procurator information comprises the formal charges and procuratorial opinions, which are utilized to instantiate the procurator agent in criminal courts.
Chinese Legal Articles. Legal articles are authored by legal experts to explain legal concepts and address issues in legal practice.
  • Civil Trial Practice Q&A literature is a curated collection of legal articles on civil trials practice, each offering an in-depth analysis of a specific legal issue. Leveraging GPT-4o, we identify character roles with relevant occupations that are likely to engage in the discussion on these analyses, and construct a list of topics grounded in these legal issues, including topics and their corresponding answers. The topics are then summarized into a shared theme. The list of topics, the theme, and associated occupations are assigned to the non-legal role agent in Knowledge Questioning.
  • HUALV website features frontier legal articles authored by legal experts in response to legal issues and user inquiries, where the articles are titled with a specific question and provide a detailed description of the background and explanation within the context. We leverage GPT-4o to divide the article into two parts: the background part, and the explanation part. The explanation part is subsequently utilized to construct coherent topic lists and the corresponding answers. The list of topics, along with background, is mapped to the non-legal role agent in Legal Consultation.
In addition, to ensure diversity, we categorize topics into two categories: Binary(i.e., yes/no) topics and open-ended questions, where responses are expected to be either statutory provisions or legal phrases. Comprehensive statistics related to these topics are provided in Sec. C. The construction prompts for the topic list are illustrated in Figure 10 and Figure 11.
Each element within these categories is systematically allocated to the appropriate agents defined in our framework.
B.2 Personality Theory Construction
Real-world clients exhibit behavioral features according to their occupation and experiences. To ensure legal environments’ better alignment with authentic legal services, we construct non-legal roles characteristics based on the Big Five Personality Traits theory, including five dimensions: extraversion, emotional stability, openness, agreeableness, and conscientiousness. Existing researchYue et al. [2025a], Chen et al. [2024] has demonstrated LLM-driven agents’ capability to follow the preset traits and role-play characters, paving way for our implementation. Specifically, we divide each dimension of Big Five Personality traits theory into three levels(high, medium, and low), and leverage GPT-4o to assign them to non-legal role agents based on their individual information. To be specific, for the non-legal role agent in Knowledge Questioning, we exploit occupations as individual information. For Legal Consultation, we utilize background to ascertain the non-legal role agent’s behavioral style. As for the Complaint Drafting, the case details are employed for behavioral styles construction. For Defence Drafting, we base the behavior modeling on the defendant’s statement of defence. To enhance the distinctness of the agents’ portrayals, we randomly reassign the dimensions with the medium level to three levels with the distribution ratio of 2:1:2. We then map five-dimensional traits into behavioral descriptions that feature role agents’ talking style leveraging GPT-4o to model role agents’ behavior.
B.3 Role Profile Configuration
Based on the complexity of the environments, three levels of roles with different granularity are configured with legal elements.
Type I: General Public represents clients of various professions and backgrounds with little legal knowledge, who seek legal counseling on a particular theme(Env.1) or a specific case(Env.2). For the group in Env.1, we assign a topic list as well as the shared theme to the role agent, and equip them with talking styles generated by the personality traits in B.2. For profiles of client agents, check Figure 13 to Figure 14.
Type II: Specific Characters are modeled with more granular information to satisfy the increased complexity of environments in Level II, including two types of settings:
  • Personal attributes feature personal traits of the characters. We include name, gender, ethnicity, address, and birth date if the character is a person. If the character is a corporation, we prompt the non-legal role agent to role-play the representative of the corporation, and equip it with the name and the address of the corporation, as well as the name of the representative. We also include talking styles to model their behaviors.
  • Case-related details depict the actions and statements of parties involved in the case. For plaintiffs, we assign claims and their evidences. For defendants, we include statement of defence and defendants’ evidences.
Level Level I Level II Level III
Env. K Q L C C D D D C I C R
98 62 93 93 93 69
Source Legal Article Civil Judgment Document Criminal Judgement Document
160 93 69
Table 4: Data statistics of J1-E VAL . Where KQ, LC, CD, DD, CI, and CR denote Knowledge Questioning, Legal Consultation, Complaint Drafting, Defence Drafting, Civil Court, and Criminal Court, respectively.
These information are assigned to the plaintiff agent(check Figure 16 and Figure 17), and the defendant agent(check Figure 18 and Figure 19).
Type III: Multi-specific Characters simulate multiple specific parties in the courtroom with the finest-grained settings and all information necessary to the case. To be specific:
  • Lawyers attend in civil courts and criminal courts, and are assigned information about their clients(i.e., the plaintiff and the defendant), as well as case-related details, including evidences, defence against accusations, and claims.
  • Defendant participates in the criminal court and is set with personal information, case details, and attitudes towards charges.
  • Procurator engages in the criminal court session, with charges as well as opinions toward the defendant.
In the Civil Court, the plaintiff lawyer and the defendant lawyer are prompted with the above configuration, as illustrated from Figure 20 to Figure 23. In addition, check Figure 25 to Figure 28 for profile settings in criminal courts.
Appendix C Data Statistics
As shown in Table 4, we meticulously select 322 quality data in J1-E val , with 160 pieces of legal articles, 93 civil judgment documents, and 69 criminal judgment documents. Legal articles are then utilized for agent configurations in Knowledge Questioning and Legal Consultation, with 98 pieces for Knowledge Questioning and 62 pieces for Legal Consultation. Civil judgment documents are utilized for Complaint & Defence Drafting, as well as Civil Court. Criminal judgment documents are employed for Criminal Court setups. Therefore, we include 508 pieces of data in J1-E val .
In addition, for topic lists in Knowledge Questioning, 286 binary topics and 188 open-ended topics are designed, with an average of 4.84 topics for each case. For topic lists in Legal Consultation, we construct 116 binary topics and 107 open-ended topics, with an average of 3.60 topics for each of the cases. The balanced percentage and diversity of different topic types have ensured a comprehensive evaluation of the model-driven agent’s performance.
Appendix D Evaluation Metric
We construct a comprehensive evaluation framework that assesses the performance of the legal agent from multiple perspectives.
D.1 Level-I
We leverage an LLM-based check mechanism to label each interaction round with its corresponding topic, which facilitates <question-answer> matching and scoring. As illustrated in Sec.B.1, the topics are divided into binary type and open-ended type, which are evaluated for Binary accuracy and Non-binary score respectively. To be specific:
  • Binary accuracy measures accuracy in binary questions with the matching method. Comparing the yes/no answer summarized by the check mechanism and the ground truth, we obtain binary accuracy: 1 for correct, 0 for incorrect, and 0.5 for conditionally qualified answers.
  • Non-binary score evaluates other questions based on the original responses of the legal agent. If the ground truth is legal provisions, we measure the score by matching method: 10 if the legal provisions are cited by the legal agent, and 0 if not. For ground truths that involve legal phrases, we leverage GPT-4o to generate a score ranging from 0 to 10. The evaluation prompt is shown in Figure 32.
Agent KQ LC CD DD CI CR
B IN N BIN B IN N BIN F OR D OC F OR D OC P FS J UD R EA L AW P FS C RI J UD R EA L AW
GPT-4o 61.46 59.08 51.14 41.60 69.46 80.57 41.72 69.09 58.23 27.31 49.14 21.11 28.08 85.51 16.46 65.51 44.45
Deepseek-v3 671B 51.00 54.53 49.60 44.81 96.34 72.49 36.77 55.09 57.85 26.67 40.11 20.14 30.34 85.51 17.77 60.14 31.01
Qwen3-it. 14B 55.42 57.94 54.03 40.20 57.85 80.94 34.84 60.98 7.15 3.76 5.81 4.34 28.26 82.61 24.10 68.41 49.84
Gemma3-it. 12B 54.49 42.89 40.93 20.11 58.49 72.79 37.85 56.05 37.87 17.31 31.72 9.74 18.38 81.16 20.75 50.29 18.84
Chatlaw2 7B 53.27 51.05 41.13 23.04 12.69 10.42 0.43 0.36 - 1.08 0.65 - - 7.25 95.08 4.64 0.91
Table 5: Performances among LLM-driven legal agents on J1-E NVS driven by Qwen3-Instruct-32B.
D.2 Level-II
We evaluate the legal document drafted by the legal agent from two perspective.
(1) Format-following score measures the correspondence of the generated document’s format with the given template. We break down the format into a set of labels $L=[L_{1},L_{2},...,L_{k}]$, which are organized in a specific sequential order $S$. As shown in Figure 29, the complaint format consists of labels of Plaintiff, Defendant, Claims, Facts and Legal Grounds, and Evidences and Sources, Names and Addresses of Witness, following a specific sequential order. As for the defence format, check Figure 30, which includes Defendant, defence arguments, case id, parties and cause of action, Evidences and Sources, names and addresses of witnesses. Through matching, we check whether the generated sequential order $S_{agent}$ follows the given order $S$, and score the accuracy of labels by comparing the generated labels $L_{agent}$ with $L$. The final result is calculated with the following equation.
$FOR=Seq*Label$ (1)
Where: $Seq$ is 1 if $S_{agent}$ is identical with $S$, and 0 if not; $Label$ is the accuracy of labels $S_{agent}$.
(2) Document score evaluates the quality of obtained information. To be specific, personal information, including plainiff’s information and the defendant’s information (i.e., name, ethnicity, address, birth date, gender), is measured through exact match(1 or 0). As for the other components(i.e., Claims, Evidences, Statement of Defence, Case Details) are evaluated with prompts in Figure 33.
Figure 8: Turn Constraints impact on Civil Court Performance
Figure 9: Turn Constraints impact on Criminal Court Performance
D.3 Level-III
We evaluate the legal agents’ performance in Civil Court and Criminal Court from two dimensions.
(1) Procedural-following score provides assessments of completeness of procedural stages, each with several mandatory actions. Only when all mandatory actions are executed is a stage considered complete. Specifically, Civil Court is divided into five stages, i.e., court preparation, court investigation, court debate, court mitigation, and court decision. As illustrated in Figure 34 and Figure 35, we evaluate the court preparation stage, court investigation stage, court debate stage, and the court mitigation stage. On the other hand, as illustrated in Figure 36, Criminal Court is divided into three stages: court preparation, court investigation, and court decision. In addition, we notice that some legal agents are prone to unilateral speech, with all the stages executed within one turn. Such performance is deemed as failure in our settings. Through matching, we calculate the completeness of procedural stages and mandatory actions, and use their average as the $PFS$ score.
(2) Judgment score evaluates the quality of the legal agent’s judgment, which encompasses two parts: the inference part and the judgment part. The inference part contains the court’s inference for the judgments as well as the applied laws. We compare the laws applied by the legal agent with the ground truth to determine the correctness, and leverage GPT-4o to evaluate the inference process. The judgment part in Civil Court includes the agent’s ruling over the case, which is evaluated with GPT-4o. The evaluation prompt could be seen in Figure 33. As for the Criminal Court, the judgment part is divided into three parts, where the criminal charge is scored through matching, and the accuracy of predicted fines and sentences is calculated with log deviation with the following equation. The average of the fine deviation and the sentence deviation is calculated as the final $VER$.
$v=|log(l_{p}+1)-log(l_{a}+1)|$ (2)
Where: $v$ is the log deviation; $l_{p}$ is the agent predicted result; and $l_{a}$ is the ground truth. Higher $v$ indicates greater deviation and thus worse performance of the legal agent.
Agent Complaint Drafting
D OC
P LA D EF C LA E VI F &R
Multilingual LLMs(close source)
GPT-4o 92.69 96.27 97.31 77.74 85.27
Claude-3.7 92.04 93.12 96.24 85.16 79.57
Multilingual LLMs(open source)
Deepseek-v3 671B 71.40 87.67 67.31 35.16 56.45
Deepseek-r1 671B 75.34 71.04 44.30 11.72 29.78
Llama3.3-it. 70B 78.85 93.12 97.96 66.67 86.02
Qwen3-it. 32B 73.98 94.27 98.49 76.02 79.35
Gemma3-it. 27B 86.52 94.41 98.39 74.73 70.11
Qwen3-it. 14B 93.98 94.48 97.10 79.68 80.54
Gemma3-it. 12B 80.43 91.47 96.88 73.76 78.17
GLM4 9B 80.80 83.87 89.35 73.55 83.76
Qwen3-it. 8B 76.27 77.13 80.97 60.86 68.28
InternLM3-it. 8B 77.42 52.11 79.57 76.99 74.84
Ministral-it. 8B - 14.19 - - -
Qwen2.5-it. 7B 84.16 94.05 91.83 52.26 84.73
Qwen3-it. 4B 89.32 91.11 94.84 76.13 66.99
Legal-specific LLMs
LawLLM 13B 0.22 2.08 4.95 3.98 3.01
Chatlaw2 7B 6.88 16.70 23.33 14.95 23.01
Table 6: Level II: DOC in Complaint Drafting. where PLA, DEF, CLA, EVI and F&R denote defendant information, claims, evidence, and fact & reason respectively.
MODEL Defence Drafting
D OC
R ES D EF E VI
Multilingual LLMs(close source)
GPT-4o 92.54 86.56 75.27
Claude-3.7 79.93 85.27 41.08
Multilingual LLMs(open source)
Deepseek-v3 671B 73.91 74.30 80.54
Deepseek-r1 671B 73.55 53.66 16.45
Llama3.3-it. 70B 71.04 71.94 44.09
Qwen3-it. 32B 73.41 79.25 37.74
Gemma3-it. 27B 78.92 80.22 47.31
Qwen3-it. 14B 77.13 76.24 33.76
Gemma3-it. 12B 68.32 83.66 39.57
GLM4 9B 69.82 61.61 25.16
Qwen3-it. 8B 25.66 27.85 14.95
InternLM3-it. 8B 47.38 70.86 43.23
Ministral-it. 8B 14.12 20.22 7.63
Qwen2.5-it. 7B 69.25 73.55 36.34
Qwen3-it. 4B 68.83 56.13 29.68
Legal-specific LLMs
LawLLM 13B - - -
Chatlaw2 7B 1.07 1.61 -
Table 7: Level II: DOC in Defence, where RES, DEF and EVI denote respondent, defence claims, evidence respectively.
Appendix E Implementation Details.
For adapted LLMs, we speed up inference using vllm Kwon et al. [2023]. Greedy decoding was used across the evaluations. For models smaller than 32B, we run evaluations using 8 * RTX 4090 GPUs with 24GB memory each. For models larger than 32B, we run evaluations using 8 * A100 GPUs with 80GB memory. For some metrics calculating in evaluation, we use GPT-4o 333gpt-40-2024-11-20. To prevent infinite interaction loops caused by limited model capabilities, we set maximum interaction rounds for each scenario: 15 for knowledge 1 questioning, 10 for legal consultation, 20 for complaint drafting, 15 for defence drafting, 50 for civil, and 35 for criminal court.
Appendix F Additional Experiment
F.1 Performance among Various Legal Agents
We provide additional results of the overall performance among LLM-driven legal agents in J1-E NVS .
  • Different J1-E NVS . Table 5 reports additional results of legal agents in J1-E NVS driven by Qwen3-Instruct-32B.
  • Complaint & Defence Drafting. Table 6 and Table 7 reports additional DOC result in Complaint Drafting and Defence Drafting. Both general and legal-specific models perform relatively well on five sub-metrics of DOC, indicating powerful information collection ability through the multi-turn conversation. However, agents driven by legal-specific or smaller models, such as LawLLM, encounter difficulty in accomplishing the assigned tasks, requiring further training.
  • Civil Court. Table 8 reports additional PFS result in Civil Court. GPT-4o and Claude-3.7 demonstrate remarkable performances in STA and ACT, indicating strong procedural-following ability of these models, while legal-specific models, the inference model, and smaller models are unable to follow civil court procedures.
  • Criminal Court. Table 9 reports additional PFS and DOC result in Criminal Court. GPT-4o, Claude-3.7, and Deepseek-v3 score 100 in UNI. In contrast, legal-specific models fail to navigate the criminal court, resulting lower UNI and poorer performance.
Agent Civil Court
P FS
S TA A CT U NI
Multilingual LLMs(close source)
GPT-4o 61.02 67.41 90.32
Claude-3.7 54.03 65.34 74.19
Multilingual LLMs(open source)
Deepseek-v3 671B 25.27 32.42 39.78
Deepseek-r1 671B 12.63 16.21 19.35
Llama3.3-it. 70B 41.13 70.89 96.77
Qwen3-it. 32B 39.78 52.19 70.97
Gemma3-it. 27B 42.47 61.62 84.95
Qwen3-it. 14B 0.81 1.99 3.23
Gemma3-it. 12B 37.10 50.79 67.74
GLM4 9B 40.32 49.38 54.84
Qwen3-it. 8B 10.22 15.47 39.78
InternLM3-it. 8B 7.26 9.35 54.84
Ministral-it. 8B 20.43 30.02 36.56
Qwen2.5-it. 7B 20.16 26.63 31.18
Qwen3-it. 4B 24.19 29.28 44.09
Legal LLMs
LawLLM 13B - - -
Chatlaw2 7B - - 5.38
Table 8: Level III: PFS in Civil Court, where STA and ACT denote stages and actions respectively, and UNI is the binary indicator.
      Agent Criminal Court
PFS JUD
STA ACT UNI SEN FINE
Multilingual LLMs (closed source)
GPT-4o 1.45 56.88 100.00 8.84 6.46
Claude-3.7 - 50.72 97.10 20.20 8.36
Multilingual LLMs (open source)
Deepseek-v3 671B - 32.79 82.61 32.52 29.01
Deepseek-r1 671B 0.72 4.17 7.25 98.53 97.24
Llama3.3-it. 70B - 43.66 91.30 15.84 23.67
Qwen3-it. 32B - 38.04 88.41 25.16 16.71
Gemma3-it. 27B 3.62 53.80 98.55 16.71 17.38
Qwen3-it. 14B 4.35 59.42 88.41 17.11 16.80
Gemma3-it. 12B 0.72 40.04 98.55 15.70 14.63
GLM4 9B 2.17 47.46 71.01 43.97 4.21
Qwen3-it. 8B - 17.21 63.77 42.52 42.48
InternLM3-it. 8B - 35.33 72.46 54.56 63.03
Ministral-it. 8B 0.72 3.62 5.80 96.30 95.83
Qwen2.5-it. 7B 6.52 36.78 59.42 45.74 48.05
Qwen3-it. 4B 0.72 30.43 71.01 40.19 36.44
Legal LLMs
LawLLM 13B - - - - -
Chatlaw2 7B - - 1.45 98.55 98.55
Table 9: Level III: PFS in Criminal Court, where STA, ACT, SEN, and FINE denote stages, actions, sentences, and fines respectively, and UNI is the binary indicator.
F.2 Analysis of Dialogue Turn Constraints on Simulated Courts Performance
We study the impact of maximum dialogue turn constraint on the performance of legal agent in civil court and criminal court driven by Qwen3-Instruct-32B. For civil court, three groups are introduced, where maximum dialogue turn is set to be 35, 40, and 50. For criminal court, we examine legal agent performance under the maximum turn constraint of 30 and 35 turns. As illustrated in Figure 8, the performance of GPT-4o in civil court demonstrates a positive correlation with the increase in dialogue turn constraint, while Deepseek-v3 and Gemma3-Instruct-12B exhibit a decreasing trend with the increase in maximum dialogue turns. The gap in performance could be attributed to the difference in long-range memory, with GPT-4o effectively retains context and executes procedural operations in detail with prolonged dialogue, while Deepseek-v3 and Gemma3-Instruct-12B suffers to retain early dialogue information, subsequently leading to diminished reasoning and judgment prediction performance. For criminal court, as shown in Figure 9, across all evaluated agents, their performance remains largely stable with the increasing dialogue constraint, as most of cases are concluded within 30 turns. The results demonstrate the value of long-range memory for legal agent performance in complex and multi-turn environment settings, which highlights the necessity for future research on a more robust memory mechanism for legal agents.
Model KQ LC CD DD CI CR
Multilingual LLMs (close source)
GPT-4o 5.91 4.92 6.70 5.17 35.30 15.29
Claude-3.7 5.89 5.04 12.77 9.18 37.48 11.61
Multilingual LLMs (open source)
Deepseek-v3 671B 6.32 4.94 12.96 4.90 6.90 7.22
Deepseek-r1 671B 5.97 5.31 10.84 4.83 3.22 1.33
Llama3.3-it. 70B 5.93 5.18 6.97 5.01 17.45 8.88
Qwen3-it. 32B 5.92 4.93 9.06 7.55 33.26 21.00
Gemma3-it. 27B 5.89 5.09 7.70 6.11 31.52 11.39
Qwen3-it. 14B 6.01 4.97 6.09 6.66 32.94 16.88
Gemma3-it. 12B 5.91 5.18 8.73 7.12 26.53 14.48
GLM4 9B 6.01 5.19 13.47 5.03 5.54 7.32
Qwen3-it. 8B 6.06 4.83 14.26 14.06 32.25 13.07
InternLM3-It. 8B 5.94 5.61 2.71 4.29 1.05 1.38
Ministral-it. 8B 6.24 6.05 1.00 2.35 29.44 27.07
Qwen2.5-it. 7B 5.93 4.93 5.60 5.24 32.35 20.00
Qwen3-it. 4B 6.02 5.17 8.12 8.18 24.15 15.52
Legal-specific LLMs
LawLLM 13B 6.14 5.42 20.00 15.00 2.45 25.30
Chatlaw2 7B 7.15 7.90 6.74 6.34 21.77 15.17
Table 10: Average turn of interaction among LLM-driven legal agents on six environments.
F.3 Analysis of interaction turn.
Table 10 presents the average number of interactions observed across the six environments. Conversely, agents that fail to complete the task tend to have few turns or directly reach the maximum turn limit. For instance, ChatGLM, InternLM3, and LawLLM achieves only 5.54, 1.05, and 2.45 turns on average in C I . The results indicate that weak interactive abilities prevent the models from sustaining focused, purposeful dialogues, often leading to off-topic, incomplete, or procedurally flawed interactions.
F.4 Analysis of confidence intervals.
To quantify the variance in outcomes across multiple runs, we conduct additional experiments along two dimensions:
  • Across different models. We select three representative models(i.e., GPT-4o, Deepseek-v3 and Gemma3-Instruct-12B), and re-run each model three times under the same settings of the main experiment in both the CD and DD scenarios. As shown in Table 13, performances of these models remain consistent across three runs, demonstrating the robustness of our framework.
  • Across different scenarios. In addition, we extend our analysis to evaluate the legal agent’s performance in other scenarios. Due to time constraints, we conduct experiments on 50% of the data samples and focus on four scenarios. Three independent re-runs are performed for GPT-4o driven legal agents. As shown in Table 14, the outcomes are stable across these runs, further proving the robustness of our framework.
Appendix G Human Evaluation
We perform a comprehensive analysis of the behavioral consistency of environment roles during interactions with three legal agents, leveraging GPT-4o and human evaluators to rate the alignment on a 10-point scale. GPT-4o assessment is performed across all samples with the prompt shown in Figure 31, while human evaluators assess 10% of the samples per scenario. As shown in Table 11 and Table 12, in interactions with various legal agents(GPT-4o, Deepseek-v3, and Qwen3-Instruct-14B), the performance of all environmental roles remains consistently high and stable. The experiment demonstrates the reliability and effectiveness of environmental roles, establishing a solid basis for the subsequent assessment of the legal agent’s performance.
Agent KQ LC CD DD CI CR
General Public General Public Plaintiff Defendant Lawyer Defendant Defendant Lawyer Procurator
Env.GPT-4o
GPT-4o 86.84 80.53 90.11 86.67 98.60 97.93 100.00 100.00 99.73
Deepseek-v3 671B 87.55 85.74 66.24 82.04 96.56 96.27 100.00 97.25 98.13
Qwen3-it. 14B 85.10 81.47 85.16 85.16 92.59 92.50 96.96 88.00 98.00
Env.Qwen3-it. 32B
GPT-4o 92.55 90.32 92.80 89.14 97.69 96.24 100.00 94.93 99.40
Deepseek-v3 671B 94.39 89.35 86.24 88.92 96.56 97.65 99.28 95.11 99.09
Qwen3-it. 14B 92.87 90.81 89.35 89.35 94.24 100.00 95.51 97.14 100.00
Table 11: Behavior consistency of various environment roles under GPT evaluation.
Agent KQ LC CD DD CI CR
General Public General Public Plaintiff Defendant Plaintiff Defendant Defendant Lawyer Procurator
Env.GPT-4o
GPT-4o 82.50 86.50 85.50 85.50 86.50 87.00 88.00 89.50 89.00
Deepseek-v3 671B 85.50 85.50 83.50 84.00 88.00 85.00 85.50 89.50 89.00
Qwen3-it. 14B 82.50 84.50 84.50 84.50 83.00 86.25 85.50 90.00 80.00
Env.Qwen3-it. 32B
GPT-4o 84.00 86.00 83.00 86.50 82.00 79.50 85.00 87.50 85.50
Deepseek-v3 671B 85.50 85.50 85.00 85.00 78.50 84.50 80.00 87.00 64.50
Qwen3-it. 14B 84.50 85.50 86.00 87.00 80.50 90.00 86.50 78.00 64.00
Table 12: Behavior consistency of various environment roles under Human evaluation.
CD DD
Agent Experiment F OR D OC F OR D OC A VE
First run 64.73 75.85 41.29 63.95 61.46
Second run 62.37 74.95 40.00 64.27 60.40
GPT-4o Third run 62.37 75.21 40.00 64.91 60.62
[…]
Second run 94.41 74.49 44.52 52.45 66.47
Deepseek-v3 671B Third run 94.41 73.71 44.52 56.36 67.25
[…]
Second run 56.13 69.00 37.85 55.81 54.70
Gemma12B-it. 12B Third run 56.13 68.87 37.85 55.42 54.57
Table 13: Performances of different legal agents in the CD and DD across three re-runs
Experiment KQ LC CD DD A VE
B IN N BIN B IN N BIN F OR D OC F OR D OC
First run 58.33 68.27 71.94 44.50 64.73 75.85 41.29 63.95 61.11
Second run 59.19 65.66 76.67 42.36 62.37 74.95 40.00 64.27 60.68
Third run 58.55 70.68 73.61 41.64 62.37 75.21 40.00 64.91 60.87
Table 14: Performance of GPT-4o in the KQ, LC, CD, & DD across three re-runs
Appendix H Prompts
We list prompts used in different environments in Table 15. In each prompt, xx needs to fill with corresponding external inputs. We meticulously design prompts for data construction and agent settings to ensure clarity and functionality. These prompts are adaptable to most LLMs, enabling the reliable configuration of environment characters.
Prompt Environment Function
Figure 10 & 11 KQ & LC Construction of Topic Lists
Figure 12 & 13 KQ General Public Prompt in Knowledge Questioning
Figure 14 & 15 LC General Public Prompt in Legal Consultation
Figure 16 & 17 CD Plaintiff Prompt in Complaint Drafting
Figure 18 & 19 DD Defendant Prompt in Defence Drafting
Figure 20, 21, 22 and 23 CI Plaintiff & Defendant Prompt in Civil Prediction
Figure 24, 25, 26, 27, and 28 CR Defendant’s Lawyer, Procurator, and Defendant Prompt in Criminal Court
Figure 29 & 30 CD & DD Complaint Format & Defence Format
Figure 31 KQ, LC, CD, & DD GPT Prompt for Consistency Evaluation
Figure 32 KQ & LC GPT Prompt for Open-ended Topics Evaluation
Figure 33 CD, DD, CI, & CR GPT Prompt for Comparative Evaluation
Figure 34 & 35 CI Detailed Court Stages in Civil Court
Figure 36 CR Detailed Court Stages in Criminal Court
Table 15: Prompts of different environments and the corresponding function.
Figure 10: Prompt of topic list construction(English Version)
Figure 11: Prompt of topic list construction(Chinese Version)
Figure 12: Prompt of Client in Knowledge Questioning(English Version)
Figure 13: Prompt of Client in Knowledge Questioning(Chinese Version)
Figure 14: Prompt of Client in Legal Consultation(English Version)
Figure 15: Prompt of Client in Legal Consultation(Chinese Version)
Figure 16: Prompt of Client in Complaint Drafting(English Version)
Figure 17: Prompt of Client in Complaint Drafting(Chinese Version)
Figure 18: Prompt of Client in Defence Drafting(English Version)
Figure 19: Prompt of Client in Defence Drafting(Chinese Version)
Figure 20: Prompt of plaintiff in Civil Court(English Version)
Figure 21: Prompt of plaintiff in Civil Court(Chinese Version)
Figure 22: Prompt of defendant in Civil Court(English Version)
Figure 23: Prompt of defendant in Civil Court(Chinese Version)
Figure 24: Prompt of lawyer in Criminal Court(English Version)
Figure 25: Prompt of lawyer in Criminal Court(Chinese Version)
Figure 26: Prompt of procurator in Criminal Court(English Version)
Figure 27: Prompt of procurator in Criminal Court(Chinese Version)
Figure 28: Prompt of defendant in Criminal Court
Figure 29: Complaint Format
Figure 30: Defence Format
Figure 31: Prompt for evaluating consistency between the Character and Statements
Figure 32: Prompt for open-ended topics evaluation
Figure 33: Prompt for comparative evaluation
Figure 34: Civil Court Stages and Actions(English Version)
Figure 35: Civil Court Stages and Actions(Chinese Version)
Figure 36: Criminal Court Stages and Actions

### v2

@@
 LawBench (EMNLP 2024) Fei et al. (2024) NLP Task Rule
-Law-Eval (DASSFA 2024) Yue et al. (2024) NLP Task Rule
+LexEval (NeurIPS 2024) Li et al. (2024) NLP Task Rule
 Law-Eval (DASSFA 2024) Yue et al. (2024) NLP Task Rule & LLM
@@
 Studies on LLM-based agents in legal contexts should focus on their ability to operate in dynamic environments, mirroring the real-world interplay of legal practitioners with evolving information and procedures.
-Ideal environment, e.g., court trials, relies on human participants with defined roles to ensure reliability, yet such setups are prohibitively costly and difficult to scale.
+The ideal environment, e.g., court trials, relies on human participants with defined roles to ensure reliability, yet such setups are prohibitively costly and difficult to scale.
 Thus, developing such a system faces two key challenges: 1) Comprehensive scenarios.
@@
 Each scenario involves diverse roles with distinct behaviors, backgrounds, and legal perspectives, all requiring a realistic and consistent portrayal.
-Beyond individuals, interaction must be dynamic yet procedurally grounded.
-To this end, we present J1-E NVS , an open-ended legal environment where agents engage in diverse legal scenarios, completing tasks through interaction with varied participants under procedural norms.
+Beyond individuals, the interaction must be dynamic and yet procedurally grounded.
+To this end, we present J1-E NVS , an open-ended legal environment where agents engage in diverse legal scenarios, completing tasks through interaction with various participants under procedural norms.
 Guided by legal experts, it curates six representative environments, categorized into three levels according to environmental complexity (role diversity, interaction demands, and procedural difficulty) as shown in Figure 1.
@@
 Each environmental role is grounded in real-world legal contexts and exhibits a distinct behavioral style, ensuring diversity and authenticity.
-Furthermore, we introduce J1-E VAL to evaluate agent capabilities within our constructed environments.
+Furthermore, we introduce J1-E VAL to evaluate the agent’s capabilities within our constructed environments.
 We design task-specific and fine-grained metrics to assess task completion and procedural compliance, providing ground truth for each task and adopting rule-based or LLM-based methods.
@@
 Our J1-E NVS covers six of the most representative real-world scenarios in Chinese legal practice.
-(B) We introduce J1-E VAL to perform fine-grained evaluation of agents across different levels of legal proficiency, ranging from trainee to lawyer to judge.
-(C) Experimental results reveal gaps between diverse agents and real-world legal demands, offering valuable insights to guide future progress.
+(B) We introduce J1-E VAL to perform fine-grained evaluations of agents across different levels of legal proficiency, ranging from trainee to lawyer to judge.
+(C) Experimental results reveal the gaps between diverse agents and real-world legal demands, offering valuable insights to guide future progress.
 (D) This work introduces a new paradigm for legal intelligence, shifting from static to dynamic.
-Beyond evaluation, it can further be extended for data generation and reinforcement learning training.
+Beyond evaluation, it can be extended to data generation and reinforcement learning training.
 Figure 2: Environment construction pipeline.
@@
 To simulate realistic scenarios, we assign different legal elements of the same event to various roles, drawing from two sources: (1) Chinese judgment documents.
-For scenarios in Level II & III, Chinese court judgments encompass all the essential components of a case’s entire procedural process, e.g., parties’ details, legal disputes, facts, rulings, and laws, which align perfectly with our requirements.
+For scenarios in Level II & III, Chinese court judgments encompass all the essential components of a case’s entire procedural process, e.g., parties’ details, legal disputes, facts, rulings, and laws, which perfectly align with our requirements.
 Specifically, we collect civil and criminal cases from judgment documents and then employ LLM to extract key elements, such as plaintiff information, claims, case details, and relevant legal provisions.
@@
 For scenarios in Level I focused on progressive questioning logic and authenticity, we collect expert-drafted legal articles from the HUALV website and the Civil Trial Practice Q&A literature.
-As these articles are structured through stepwise questions or thematic headings, we prompt LLM to leverage this inherent structure to generate coherent topic lists and corresponding answers.
+As these articles are structured through stepwise questions or thematic headings, we prompt an LLM to leverage this inherent structure to generate coherent topic lists and corresponding answers.
 Topics are classified as binary (yes/no) and open-ended questions.
@@
 (3) Type III: Multi-specific Characters hold specific responsibilities within courtroom proceedings and are assigned real-world legal information corresponding to their roles.
-For example, the Lawyers in the civil court possess detailed case information extracted from civil judgment documents, e.g., their client information, case details, claims & defence.
+For example, Lawyers in the civil court possess detailed case information extracted from civil judgment documents, e.g., their client information, case details, claims & defence.
 In criminal court scenarios sourced from criminal judgment documents, the Procurator agent is assigned to charges and opinions, while the Defendant agent is assigned case details and a stated stance toward the charges.
@@
 Since legal agents lack information about the public, they actively ask for relevant details to provide better responses.
-Similar to KQ, the interaction continues until all question lists are resolved.
+Similarly to KQ, the interaction continues until all questions are resolved.
 Level II: Complaint Drafting (C D ) and Defence Drafting (D D ).
@@
 (2) D D .
-In response to the defendant’s need for a statement of defence, the legal agent engages in in-depth dialogue with the defendant, drawing on the complaint to organize arguments and ultimately generate the defence document.
+In response to the defendant’s need for a statement of defence, the legal agent engages in an in-depth dialogue with the defendant, drawing on the complaint to organize arguments and ultimately generate the defence document.
 The end of conditions is similar to complaint drafting.
@@
 3.1 Data Details
-Following the details in Sec.2.1, we collected legal articles and documents from the year 2025 to ensure the timeliness.
+Following the details in Sec.2.1, we collected legal articles and documents from the year 2025 to ensure timeliness.
 After systematic processing, J1-E VAL sets out 508 distinct environments, including 160 at Level I, 186 at Level II, and 192 at Level III.
-The detailed statistics are presented in Table 4 in Appendix.
+Detailed statistics are presented in Table 4 in Appendix.
 Moreover, J1-E VAL encompasses a diverse and nuanced set of legal attributes aligned with real-world legal practice.
@@
 The two setting types provide a comprehensive assessment: for binary questions, outputs are summarized as "yes" or "no," while for non-binary questions, original responses are used for evaluation.
-  • Binary accuracy (B IN ) [0, 0.5, 1] measures accuracy in binary questions by matching method: 1 for correct, 0 for incorrect, and 0.5 for conditionally qualified answers.
+  • Binary accuracy (B IN ) [0, 0.5, 1] measures accuracy in binary questions using a matching method: 1 for correct answers, 0 for incorrect, and 0.5 for conditionally qualified answers.
   • Non-binary score (NB IN ) [0-10] evaluates the response based on ground truth in terms of legal points, logic, and factual accuracy. We use GPT-4o to implement it.
 (2) Level II.
-This stage involves the task of document generation, where the quality reflects the completeness of litigant information and evidence, the understanding of claims, and the appropriate use of document templates.
+This stage involves the task of document generation, where quality reflects the completeness of litigant information and evidence, the understanding of the claims, and the appropriate use of document templates.
 Two metrics are used:
@@
 (3) Level III.
-This stage evaluates the quality of the final judgment and procedural compliance.
+This stage evaluates the quality of the final judgment and the procedural compliance.
 Given the differences between criminal and civil judgments, we adopt distinct methods.
 The evaluation metrics are as follows:
-  • Procedural-following score (P FS ) [0-1] evaluates completeness of procedural stage.
+  • Procedural-following score (P FS ) [0-1] evaluates the completeness of the procedural stage.
 The court contains several stages, each comprising a set of mandatory actions defined based on the legal professional’s setting.
-A stage is deemed complete only when all presetting actions within that stage are correctly matched.
+A stage is deemed complete only when all pre-setting actions within that stage are correctly matched.
   • Judgment score (J UD ) [0-1 or 1-10] evaluates the judgment quality.
-For civil court, LLM evaluates the judgment against the ground truth regarding legal points, logicality, and factual accuracy.
-For criminal court, we use matching to calculate the crime (C RI ) accuracy.
+For the civil court, LLM evaluates the judgment against the ground truth regarding legal points, logicality, and factual accuracy.
+For the criminal court, we use matching to calculate the crime (C RI ) accuracy.
 Additionally, following Chen et al.
@@
 Chatlaw2 7B 60.64 53.11 50.40 25.01 22.37 16.97 1.29 0.89 - 1.08 0.54 - - - 98.55 - -
-Table 2: Performances among LLM-driven legal agents on J1-E NVS , where the bolded score indicates the best performance, while the underlined score represents the second-best.
+Table 2: performance among LLM-driven legal agents on J1-E NVS , where the bolded score indicates the best performance, while the underlined score represents the second-best.
 Note that “-” denotes failure to accomplish the task, and lower V ER indicates higher performance.
@@
 4.2 Performance among Various Legal Agents
-In this section, we use GPT-4o Achiam et al. (2023) to drive our environments, where LLM-base agents enter these to accomplish set legal tasks.
+In this section, we use GPT-4o Achiam et al. (2023) to drive our environments, where LLM-based agents enter to accomplish set legal tasks.
 Overall Performance.
 To assess overall performance, we compute the average scores across all environments except V ER , and rank the models by size.
-As shown in Figure 5, GPT-4o achieved the best performance, showing strong legal intelligence.
+As shown in Figure 5, GPT-4o achieves the best performance, showing strong legal intelligence.
 Qwen3-Instruct-32B demonstrates performance beyond expectations, surpassing Deepseek series.
-Notably, although the legal-specific LLMs perform comparably to the GPT series on existing legal benchmarks Fei et al.
+Notably, although the legal-specific LLMs perform comparably to GPT series on existing legal benchmarks Fei et al.
 (2024); Yue et al.
@@
 However, their performance drops on the more interactive L C task, which requires proactive engagement.
-Notably, even legal-specific agents built on older base can match the performance of newer LLMs.
+Notably, even legal-specific agents built on an older base can match the performance of newer LLMs.
 Results indicate that current agents possess legal knowledge and can apply it effectively in basic legal consultations.
@@
 In summary, effective reasoning and adherence to procedural protocols within long legal contexts are essential for achieving accurate legal judgments.
-Detailed results of judgment components are shown in Tables 8 and 9 of Appendix.
+Detailed results of the judgment components are shown in Tables 8 and 9 of Appendix.
 Performance of Procedural-Following Ability
@@
 As shown in Figure 5, legal agents struggle to complete necessary actions in the Preparation stage in both civil and criminal courts.
-In addition, despite fewer stages and actions, navigating through the criminal court poses greater challenges for legal agents, as the environment contains more characters that increase the complexity of interaction.
+In addition, despite fewer stages and actions, navigating the criminal court poses greater challenges for legal agents, as the environment contains more characters that increase the complexity of the interaction.
 These findings underscore the need for enhanced procedural learning for legal agents, as well as the importance of strengthening their communication and coordination with different characters.
@@
 (2025) have shown the strong role-playing and community modeling abilities of LLMs.
-In this section, we use GPT-4o and Qwen3-Instruct-32B to drive the J1-E NVS separately, demonstrating the reliability.
+In this section, we use GPT-4o and Qwen3-Instruct-32B to drive the J1-E NVS separately, demonstrating reliability.
 Agent Env.GPT-4o Env.Qwen3-it. 32B
@@
 We analyze the behavioral consistency of environment roles during interactions with three legal agents.
-Both GPT-4o and human evaluators rate the consistency between each role’s profile and its behavior on a scale of 1 to 10, GPT-4o evaluated all samples, while human evaluators assessed 10% of the samples per scenario.
-The average scores are shown in the Table 3.
+Both GPT-4o and human evaluators rate the consistency between each role’s profile and its behavior on a scale of 1 to 10, GPT-4o evaluates all samples, while human evaluators assesse 10% of the samples per scenario.
+The average scores are shown in Table 3.
 Across interactions with various legal agents (GPT-4o, Deepseek-v3, and Qwen3-Instruct-14B), the ratings remain consistently high and stable.
-Furthermore, the scores for interactions with the same agent are comparable regardless of environments.
+Furthermore, the scores for interactions with the same agent are comparable regardless of the environment.
 This experiment validates the reliability and effectiveness of our environment, laying a solid foundation for assessing agents’ performance in interactive legal scenarios.
@@
 Analysis with different J1-E NVS .
-To further validate the stability, we employ a smaller, publicly accessible LLM (Qwen3-Instuct-32B) to drive our environment, and evaluate the overall performance of five legal agents, i.e, GPT-4o, Deepseek-V3, Qwen3-Instuct-14B, Gemma3-Instuct-14B and Chatlaw2.
-Figure 7 illustrates that: 1) all five agents successfully completed their respective tasks within the Qwen-based environment and achieved the same rankings as in the GPT-4o-based environment; 2) under the Qwen3-Instruct-32B environment, the performance of each agent is lower than that in the GPT-based setup, which is an acceptable outcome given the gap in model size and capability.
+To further validate stability, we employ a smaller, publicly accessible LLM (Qwen3-Instruct-32B) to drive our environment, and evaluate the overall performance of five legal agents, i.e, GPT-4o, Deepseek-V3, Qwen3-Instruct-14B, Gemma3-Instruct-14B and Chatlaw2.
+Figure 7 illustrates that: 1) all five agents successfully completed their respective tasks within the Qwen-based environment and achieves the same rankings as in the GPT-4o-based environment; 2) under the Qwen3-Instruct-32B environment, the performance of each agent is lower than that in the GPT-based setup, which is an acceptable outcome given the gap in model size and capability.
 Detailed results are shown in Table 5.
@@
 Human evaluators are asked to evaluate the performance of legal agents, and the difference between the human score and the corresponding GPT-4o score is plotted in Figure 7.
-We find that the scores provided by GPT-4o and human evaluators largely(over 76%) remain consistent.
+We find that the scores provided by GPT-4o and human evaluators largely remain consistent(over 76%).
 These results demonstrate the reliability and robustness of our evaluation framework, which enhances the validity of the findings of the legal agents.
@@
 This paper introduces J1-E NVS , an environment designed to simulate agents’ legal interactions in various legal scenarios.
-We further presents J1-E VAL to assess agents’ ability to operate effectively within dynamic environments.
+We further present J1-E VAL to assess agents’ ability to operate effectively within dynamic environments.
 Distinct from prior benchmarks, our framework emphasizes interactivity and encompasses a broad range of realistic legal tasks.
 Experimental results demonstrate the stability and effectiveness of dynamic environments in J1-E NVS .
-By evaluating a wide range of existing general-purpose and legal-specific LLMs as legal agents, the study uncovers both notable advancements and persistent challenges faced by current agents.
+By evaluating a wide range of existing general-purpose and legal-specific LLMs as legal agents, the study uncovers both notable advances and persistent challenges faced by current agents.
 Our framework not only functions as an evaluation system but also serves as a reliable environmental toolkit for data generation or reinforcement learning training.
@@
 [2020], Ji et al.
-[2021] relies on task-specific models and curated datasets for specific legal tasks.
+[2021] rely on task-specific models and curated datasets for specific legal tasks.
 With LLMs’ strong generalizability Brown et al.
@@
 [2024].
-Specifically, Law-Eval Yue et al.
+Specifically, LegalBenchGuha et al.
+[2023] and LawBenchFei et al.
+[2024] evaluate various dimensions of legal competence by consolidating existing legal tasks into an LLM-adapted format.
+Law-Eval Yue et al.
 [2024] evaluates the LLM using multiple-choice questions from law-related exams.
-LawBenchFei et al.
-[2024]and LegalBenchGuha et al.
-[2023] evaluate various dimensions of legal competence by consolidating existing legal tasks into an LLM-adapted format.
+Inspired by Bloom’s taxonomy, LexEval Li et al.
+[2024] proposes a new taxonomy of legal cognitive abilities to organize different tasks.
 Furthermore, LAiW Dai et al.
 [2023] follows the logic of legal practice of Syllogism by dividing the legal capabilities of LLMs into three levels.
-Inspired by Bloom’s taxonomy, LexEval Li et al.
-[2024] proposes a new taxonomy of legal cognitive abilities to organize different tasks.
 Existing benchmarks adopt a static framework centered on a singular dimension of task competence.
@@
 [2025].
-Based on role agents, some works attempt to build more complex scenarios Stade et al.
+Based on role agents, some works attempts to build more complex scenarios Stade et al.
 [2024], Fan et al.
@@
 B.1 Legal Sources Processing
-We utilize two primary legal sources to construct the role-specific settings.
+We utilize two primary legal sources to construct role-specific settings.
 Chinese Judgment Documents.
-Judgment documents are official court records of judicial proceedings, outcomes, and factual details regarding involving parties.
+Judgment documents are official court records of judicial proceedings, outcomes, and factual details regarding parties involved.
 We collect both civil and criminal judgment documents, and employ an extraction model (GPT-4o) to identify and structure relevant legal elements.
@@
 For individual-specific elements, we collect name, gender, ethnicity, birthdate, and residential address.
-For legal entities(corporations), personal attributes include the corporate name, address, and the identity of the legal representative.
-These elements are subsequently assigned to the roles of plaintiff and defendant in Complaint & Defence Drafting, as well as Civil Court.
-  • Case information captures the procedural details of the case, including plaintiff claims, case details, the defendant’s defence against accusations, and evidences.
+For legal entities(corporations), personal attributes include the corporate name, address, and identity of the legal representative.
+These elements are subsequently assigned to the roles of plaintiff and defendant in Complaint & Defence Drafting, as well as the Civil Court.
+  • Case information captures the procedural details of the case, including plaintiff claims, case details, the defendant’s defence against accusations, and evidence.
   • Court Information includes all the necessary elements for court rulings, containing the final judgment, court findings about the case, and laws applied by the court in the case.
@@
   • Court information contains the final verdict, court findings, and legal provisions cited by the judge, resembling the information extracted from civil documents.
-  • Procurator information comprises the formal charges and procuratorial opinions, which are utilized to instantiate the procurator agent in criminal courts.
+  • Procurator information comprises formal charges and procuratorial opinions, which are utilized to instantiate the procurator agent in criminal courts.
 Chinese Legal Articles. Legal articles are authored by legal experts to explain legal concepts and address issues in legal practice.
@@
 [2025a], Chen et al.
-[2024] has demonstrated LLM-driven agents’ capability to follow the preset traits and role-play characters, paving way for our implementation.
+[2024] has demonstrated LLM-driven agents’ capability to follow the preset traits and role-play characters, paving the way for our implementation.
 Specifically, we divide each dimension of Big Five Personality traits theory into three levels(high, medium, and low), and leverage GPT-4o to assign them to non-legal role agents based on their individual information.
@@
 To enhance the distinctness of the agents’ portrayals, we randomly reassign the dimensions with the medium level to three levels with the distribution ratio of 2:1:2.
-We then map five-dimensional traits into behavioral descriptions that feature role agents’ talking style leveraging GPT-4o to model role agents’ behavior.
+We then map five-dimensional traits into behavioral descriptions that feature role agents’ talking styles leveraging GPT-4o to model role agents’ behavior.
 B.3 Role Profile Configuration
@@
 For profiles of client agents, check Figure 13 to Figure 14.
-Type II: Specific Characters are modeled with more granular information to satisfy the increased complexity of environments in Level II, including two types of settings:
+Type II: Specific Characters are modeled with more granular information to satisfy the increased complexity of environments at Level II, including two types of settings:
   • Personal attributes feature personal traits of the characters.
@@
 We also include talking styles to model their behaviors.
-  • Case-related details depict the actions and statements of parties involved in the case.
-For plaintiffs, we assign claims and their evidences.
-For defendants, we include statement of defence and defendants’ evidences.
+  • Case-related details depict the actions and statements of the parties involved in the case.
+For plaintiffs, we assign the claims and their evidence.
+For defendants, we include the statement of defence and defendants’ evidence.
 Level Level I Level II Level III
@@
 Where KQ, LC, CD, DD, CI, and CR denote Knowledge Questioning, Legal Consultation, Complaint Drafting, Defence Drafting, Civil Court, and Criminal Court, respectively.
-These information are assigned to the plaintiff agent(check Figure 16 and Figure 17), and the defendant agent(check Figure 18 and Figure 19).
+This information is assigned to the plaintiff agent(check Figure 16 and Figure 17), and the defendant agent(check Figure 18 and Figure 19).
 Type III: Multi-specific Characters simulate multiple specific parties in the courtroom with the finest-grained settings and all information necessary to the case. To be specific:
-  • Lawyers attend in civil courts and criminal courts, and are assigned information about their clients(i.e., the plaintiff and the defendant), as well as case-related details, including evidences, defence against accusations, and claims.
+  • Lawyers attend in civil courts and criminal courts, and are assigned information about their clients(i.e., the plaintiff and the defendant), as well as case-related details, including evidence, defence against accusations, and claims.
   • Defendant participates in the criminal court and is set with personal information, case details, and attitudes towards charges.
@@
 Legal articles are then utilized for agent configurations in Knowledge Questioning and Legal Consultation, with 98 pieces for Knowledge Questioning and 62 pieces for Legal Consultation.
-Civil judgment documents are utilized for Complaint & Defence Drafting, as well as Civil Court.
+Civil judgment documents are utilized for Complaint & Defence Drafting, as well as the Civil Court.
 Criminal judgment documents are employed for Criminal Court setups.
@@
   • Non-binary score evaluates other questions based on the original responses of the legal agent.
-If the ground truth is legal provisions, we measure the score by matching method: 10 if the legal provisions are cited by the legal agent, and 0 if not.
+If the ground truth is legal provisions, we measure the score by a matching method: 10 if the legal provisions are cited by the legal agent, and 0 if not.
 For ground truths that involve legal phrases, we leverage GPT-4o to generate a score ranging from 0 to 10.
@@
 Chatlaw2 7B 53.27 51.05 41.13 23.04 12.69 10.42 0.43 0.36 - 1.08 0.65 - - 7.25 95.08 4.64 0.91
-Table 5: Performances among LLM-driven legal agents on J1-E NVS driven by Qwen3-Instruct-32B.
+Table 5: performance among LLM-driven legal agents on J1-E NVS driven by Qwen3-Instruct-32B.
 D.2 Level-II
-We evaluate the legal document drafted by the legal agent from two perspective.
-(1) Format-following score measures the correspondence of the generated document’s format with the given template.
+We evaluate the legal document drafted by the legal agent from two perspectives.
+(1) Format-following score measures the correspondence of the generated document format with the given template.
 We break down the format into a set of labels $L=[L_{1},L_{2},...,L_{k}]$, which are organized in a specific sequential order $S$.
-As shown in Figure 29, the complaint format consists of labels of Plaintiff, Defendant, Claims, Facts and Legal Grounds, and Evidences and Sources, Names and Addresses of Witness, following a specific sequential order.
-As for the defence format, check Figure 30, which includes Defendant, defence arguments, case id, parties and cause of action, Evidences and Sources, names and addresses of witnesses.
+As shown in Figure 29, the complaint format consists of labels of Plaintiff, Defendant, Claims, Facts and Legal Grounds, and evidence and Sources, Names and Addresses of Witness, following a specific sequential order.
+As for the defence format, check Figure 30, which includes Defendant, defence arguments, case id, parties and cause of action, evidence and Sources, names and addresses of witnesses.
 Through matching, we check whether the generated sequential order $S_{agent}$ follows the given order $S$, and score the accuracy of labels by comparing the generated labels $L_{agent}$ with $L$.
@@
 (2) Document score evaluates the quality of obtained information.
-To be specific, personal information, including plainiff’s information and the defendant’s information (i.e., name, ethnicity, address, birth date, gender), is measured through exact match(1 or 0).
-As for the other components(i.e., Claims, Evidences, Statement of Defence, Case Details) are evaluated with prompts in Figure 33.
+To be specific, personal information, including the plaintiff’s information and the defendant’s information (i.e., name, ethnicity, address, birth date, gender), is measured through exact match(1 or 0).
+As for the other components(i.e., Claims, evidence, Statement of Defence, Case Details), they are evaluated with prompts in Figure 33.
 Figure 8: Turn Constraints impact on Civil Court Performance
@@
 We evaluate the legal agents’ performance in Civil Court and Criminal Court from two dimensions.
-(1) Procedural-following score provides assessments of completeness of procedural stages, each with several mandatory actions.
+(1) Procedural-following score provides assessments of the completeness of the procedural stages, each with several mandatory actions.
 Only when all mandatory actions are executed is a stage considered complete.
-Specifically, Civil Court is divided into five stages, i.e., court preparation, court investigation, court debate, court mitigation, and court decision.
+Specifically, the civil Court is divided into five stages, i.e., court preparation, court investigation, court debate, court mitigation, and court decision.
 As illustrated in Figure 34 and Figure 35, we evaluate the court preparation stage, court investigation stage, court debate stage, and the court mitigation stage.
@@
 In addition, we notice that some legal agents are prone to unilateral speech, with all the stages executed within one turn.
-Such performance is deemed as failure in our settings.
-Through matching, we calculate the completeness of procedural stages and mandatory actions, and use their average as the $PFS$ score.
+This performance is deemed a failure in our settings.
+Through matching, we calculate the completeness of the procedural stages and mandatory actions, and use their average as the $PFS$ score.
 (2) Judgment score evaluates the quality of the legal agent’s judgment, which encompasses two parts: the inference part and the judgment part.
@@
 We compare the laws applied by the legal agent with the ground truth to determine the correctness, and leverage GPT-4o to evaluate the inference process.
-The judgment part in Civil Court includes the agent’s ruling over the case, which is evaluated with GPT-4o.
+The judgment part in the Civil Court includes the agent’s ruling over the case, which is evaluated with GPT-4o.
 The evaluation prompt could be seen in Figure 33.
@@
 $v=|log(l_{p}+1)-log(l_{a}+1)|$ (2)
-Where: $v$ is the log deviation; $l_{p}$ is the agent predicted result; and $l_{a}$ is the ground truth. Higher $v$ indicates greater deviation and thus worse performance of the legal agent.
+Where: $v$ is the logarithmic deviation; $l_{p}$ is the agent predicted result; and $l_{a}$ is the ground truth. Higher $v$ indicates greater deviation and thus worse performance of the legal agent.
 Agent Complaint Drafting
@@
 For models larger than 32B, we run evaluations using 8 * A100 GPUs with 80GB memory.
-For some metrics calculating in evaluation, we use GPT-4o 333gpt-40-2024-11-20.
+For some metrics calculating in the evaluation, we use GPT-4o 333gpt-40-2024-11-20.
 To prevent infinite interaction loops caused by limited model capabilities, we set maximum interaction rounds for each scenario: 15 for knowledge 1 questioning, 10 for legal consultation, 20 for complaint drafting, 15 for defence drafting, 50 for civil, and 35 for criminal court.
@@
   • Complaint & Defence Drafting.
-Table 6 and Table 7 reports additional DOC result in Complaint Drafting and Defence Drafting.
+Table 6 and Table 7 report additional DOC results in Complaint Drafting and Defence Drafting.
 Both general and legal-specific models perform relatively well on five sub-metrics of DOC, indicating powerful information collection ability through the multi-turn conversation.
@@
   • Civil Court.
-Table 8 reports additional PFS result in Civil Court.
-GPT-4o and Claude-3.7 demonstrate remarkable performances in STA and ACT, indicating strong procedural-following ability of these models, while legal-specific models, the inference model, and smaller models are unable to follow civil court procedures.
+Table 8 reports additional PFS results in the Civil Court.
+GPT-4o and Claude-3.7 demonstrate remarkable performance in STA and ACT, indicating strong procedural-following ability of these models, while legal-specific models, the inference model, and smaller models are unable to follow civil court procedures.
   • Criminal Court.
-Table 9 reports additional PFS and DOC result in Criminal Court.
+Table 9 reports additional PFS and DOC results in the Criminal Court.
 GPT-4o, Claude-3.7, and Deepseek-v3 score 100 in UNI.
@@
 F.2 Analysis of Dialogue Turn Constraints on Simulated Courts Performance
-We study the impact of maximum dialogue turn constraint on the performance of legal agent in civil court and criminal court driven by Qwen3-Instruct-32B.
-For civil court, three groups are introduced, where maximum dialogue turn is set to be 35, 40, and 50.
+We study the impact of maximum dialogue turns constraint on the performance of the legal agent in civil court and criminal court driven by Qwen3-Instruct-32B.
+For civil court, three groups are introduced, where the maximum dialogue turns are set to 35, 40, and 50.
 For criminal court, we examine legal agent performance under the maximum turn constraint of 30 and 35 turns.
 As illustrated in Figure 8, the performance of GPT-4o in civil court demonstrates a positive correlation with the increase in dialogue turn constraint, while Deepseek-v3 and Gemma3-Instruct-12B exhibit a decreasing trend with the increase in maximum dialogue turns.
-The gap in performance could be attributed to the difference in long-range memory, with GPT-4o effectively retains context and executes procedural operations in detail with prolonged dialogue, while Deepseek-v3 and Gemma3-Instruct-12B suffers to retain early dialogue information, subsequently leading to diminished reasoning and judgment prediction performance.
-For criminal court, as shown in Figure 9, across all evaluated agents, their performance remains largely stable with the increasing dialogue constraint, as most of cases are concluded within 30 turns.
+The gap in performance could be attributed to the difference in long-range memory, with GPT-4o effectively retaining context and executing procedural operations in detail with prolonged dialogue, while Deepseek-v3 and Gemma3-Instruct-12B suffer from the retention of early dialogue information, subsequently leading to diminished reasoning and judgment prediction performance.
+For the criminal court, as shown in Figure 9, across all evaluated agents, their performance remains largely stable with the increasing dialogue constraint, as most of cases are concluded within 30 turns.
 The results demonstrate the value of long-range memory for legal agent performance in complex and multi-turn environment settings, which highlights the necessity for future research on a more robust memory mechanism for legal agents.
@@
 Conversely, agents that fail to complete the task tend to have few turns or directly reach the maximum turn limit.
-For instance, ChatGLM, InternLM3, and LawLLM achieves only 5.54, 1.05, and 2.45 turns on average in C I .
+For instance, ChatGLM, InternLM3, and LawLLM achieve only 5.54, 1.05, and 2.45 turns on average at C I .
 The results indicate that weak interactive abilities prevent the models from sustaining focused, purposeful dialogues, often leading to off-topic, incomplete, or procedurally flawed interactions.
@@
 We select three representative models(i.e., GPT-4o, Deepseek-v3 and Gemma3-Instruct-12B), and re-run each model three times under the same settings of the main experiment in both the CD and DD scenarios.
-As shown in Table 13, performances of these models remain consistent across three runs, demonstrating the robustness of our framework.
+As shown in Table 13, performance of these models remains consistent across three runs, demonstrating the robustness of our framework.
   • Across different scenarios.
@@
 Gemma12B-it. 12B Third run 56.13 68.87 37.85 55.42 54.57
-Table 13: Performances of different legal agents in the CD and DD across three re-runs
+Table 13: performance of different legal agents in the CD and DD across three re-runs
 Experiment KQ LC CD DD A VE